OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x − y = 6 4x − 2y = 12 0 4 x can be any number as there are infinitely many solutions to the system. There is no x value as there is no solution to the system.

OpenStudy (diyadiya):

2x − y = 6 4x − 2y = 12 Multiply the firrst equation by 2 what do you get?

OpenStudy (anonymous):

12

OpenStudy (diyadiya):

2(2x-y=6) 2(2x) - 2(y) =2(6) 4x - 2y =12 Now you can see that the two equations are identical.Thus when 2 equations are identical then there is infinitely many solutions

OpenStudy (anonymous):

that makes since mind helping me with more :)

OpenStudy (diyadiya):

You can post your new question :)